Scott and Lauren are making a brick path in the school garden. Scott places 6 rows of bricks with 2 bricks in each row. Lauren places 4 rows of bricks with 3 bricks in each row. Scott turns to Lauren and says , "Look! This is what we did in math. We’re using the communitive property of multiplication." Lauren disagrees. Who is right? 1.draw an arry to show how Scott played his bricks. Write a multiplication sentence for the arry. 2.draw an arry to show how Lauren placed her bricks. Write a multiplication sentence fo the arry. 3.Explain who is correct.
Answer by richard1234(7193)

This is a really silly question...

Technically Scott is wrong and Lauren is right, since the commutative property basically states that a*b = b*a, which cannot be used to show that 6*2 = 4*3 (even though it is correct).

Also, it is spelled "array."